OpenStudy (anonymous):

could someone plezzz help me ill give a medal!!!!!! Divide and then reduce to lowest terms. 5/8 ÷ 3/4

OpenStudy (anonymous):

\[\huge \frac{5}{8} \div \frac{3}{4} =\frac{5}{8} \times \frac{4}{3}=\frac{5}{2} \times \frac{1}{3}\] \[\huge \frac{5 \times 1}{2 \times 3} = \frac{ 5}{6}\] is required answer

OpenStudy (biokiiller396):

.625/.75=.83 repeating which is 5/6

OpenStudy (anonymous):

thank you :)